A dump truck is unloading sand which is falling into a conical pile. The volume of a cone is V=1/3 #r^2 h where v is the volume, r is radius and h is height of the cone. Find the height of the cone when diameter of the pile is 12 feet and the volume of the sand is 100feet^3. Round your answer to the nearest tenth.
Lee Ann

Answer by Nate(3500)

I will solve, but the volume of a cone is (1/3)(pi)(r^2)(h)
To solve for the height, you must get "h" by itself....
V=(1/3)(pi)(r^2)(h)
(V)/((1/3)(pi)(r^2))=h
Know, we plug in the numbers..... (diameter is 12, radius is 6, so r^2 is 36)
100/((1/3)(3.14)(36)=h
about 2.7 feet is your height!
the answer is (if you want to check for yourself) is....
2.6539278131 and more